# ECG Semantic Integrator (ESI): A Foundation ECG Model Pretrained with LLM-Enhanced Cardiological Text

### v1

@@ 4.4 Downstream Task: ECG-based User Identification
ECG generally shows unique patterns that can distinguish individuals, which makes them suitable for subject identification tasks and offers potential advantages over other biometric traits (Melzi et al., 2023). For example, compared to identifying persons with facial images, using ECG can further protect users’ privacy during usage. In this study, we leverage the PTB-XL (Wagner et al., 2020) and ICBEB (Liu et al., 2018) datasets, also used in the arrhythmia diagnosis task, to design a one-shot learning benchmark for ECG identification. Similar to the arrhythmia diagnosis task in Section 4.3, we focus solely on the test splits of the PTB-XL and ICBEB datasets. For PTB-XL, we select 5-second signal sequences from each of the 1907 subjects for both training and testing, which results in a 1907-class classification task. Similarly, we select 4-second samples and classes from the ICBEB dataset, which forms a 689-class classification task.
In this task, we compare the performance of the proposed ESI method with various approaches under linear probing and fine-tuning settings. For the specialized supervised methods, we compared our method with established supervised learners, including LSTM, XResNet101, ResNet50, and ensemble methods (Strodthoff et al., 2020). Besides the supervised learning methods, we also evaluated the ESI method against signal-focused SSL approaches including SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), CLOCS (Kiyasseh et al., 2021), and LEAVES (Yu et al., 2022a) under the linear probing and fine-tuning settings. Due to the patient de-identification during pretraining, we cannot perform zero-shot learning for this identification task, as the model has not been exposed to identifiable patient information.
@@
Table 4: Evaluation results of 1-shot ECG-based user identification task under settings including supervised learning, linear probing, and fine-tuning. The metric used is area under the ROC curve (AUC) and accuracy score (ACC). The best results are highlighted in bold.
@@
LSTM (Strodthoff et al., 2020) 0.905 0.446 0.916 0.608
XResNet101 (Strodthoff et al., 2020) 0.917 0.477 0.922 0.624
ResNet50 (Strodthoff et al., 2020) 0.922 0.493 0.930 0.636
Ensemble (Strodthoff et al., 2020) 0.925 0.502 0.936 0.653
ConvNeXt-Tiny (Woo et al., 2023) 0.919 0.481 0.934 0.647
ConvNeXt-Base (Woo et al., 2023) 0.920 0.488 0.932 0.640
@@
SimCLR (Chen et al., 2020) 0.802 0.174 0.842 0.254
BYOL (Grill et al., 2020) 0.835 0.236 0.857 0.281
CLOCS (Kiyasseh et al., 2021) 0.814 0.199 0.835 0.237
LEAVES (Yu et al., 2022a) 0.841 0.248 0.855 0.276
(Ours) ESI-tiny 0.923 0.511 0.937 0.654
(Ours) ESI 0.926 0.518 0.942 0.663
@@
SimCLR (Chen et al., 2020) 0.934 0.522 0.945 0.673
BYOL (Grill et al., 2020) 0.942 0.547 0.951 0.686
CLOCS (Kiyasseh et al., 2021) 0.929 0.516 0.940 0.666
LEAVES (Yu et al., 2022a) 0.944 0.550 0.953 0.688
(Ours) ESI-tiny 0.966 0.591 0.980 0.747
(Ours) ESI 0.970 0.608 0.985 0.762
@@ 5.2 Ablation Study: Selection of Backbone for Signal Encoder
Table 5b(b) summarizes the results. ConvNeXt-based models (ConvNeXt-tiny and ConvNeXt-base) achieved the highest AUC scores on both arrhythmia diagnosis and ECG identification tasks. Notably, ConvNeXt-base, the largest model with the most parameters (85.56M), provides the best overall performance (AUC of 0.932 for arrhythmia diagnosis and 0.926 for ECG identification). ConvNeXt-tiny, a more parameter-efficient option (26.81M), shows a slightly lower performance compared to the base version, but also achieves competitive AUC scores with a significantly lower parameter footprint. With the same-level size as ConvNeXt models in both tasks, the ViT model shows substantially lower performances, which might indicate the convolutional kernel could be more suitable in processing ECG signals. Moreover, the XResNet1D101 backbone displays lower AUC scores compared to ConvNeXt architectures. This can due to the significantly smaller parameter size of the XResNet model.
@@ 5.3 Ablation Study: Impact of Pretraining Data Sizes and Distribution Shifts
In this ablation study, we also investigate the impact of pretraining dataset size on the performance of our ESI-tiny model in two arrhythmia classification datasets including PTB-XL and ICBEB. We artificially change the size of the unlabeled pretraining dataset from 0 to all available samples by randomly sampling and selecting from all the pretraining samples. After ESI-tiny encoders are trained, the models are then evaluated on both datasets using a linear probing approach with frozen encoder weights, and the AUC score is employed as the performance metric.
@@ 6 Conclusion
This study introduces a novel multimodal contrastive pretraining framework to enhance the quality and robustness of representations learned from ECG signals. To address the lack of descriptive text associated with ECGs, we propose a retrieval-augmented generation (RAG) pipeline called the Cardio Query Assistant (CQA). This pipeline generates detailed textual descriptions for ECG data with demographic information, potential conditions, and waveform patterns. Inspired by the success of multimodal pretraining strategies in vision-language tasks, we develop the ECG Semantic Integrator (ESI). This framework integrates both contrastive and captioning capabilities to foster a deeper semantic understanding of ECG signals. Our evaluation validates the effectiveness of the proposed approach across several downstream tasks. The ESI method demonstrates improvement in arrhythmia diagnosis and ECG-based user identification tasks by outperforming strong baselines that cover supervised learning and SSL approaches. These results with the ablation studies highlight the benefits of multimodal learning for ECG analysis and the value of integrating captioning loss with contrastive pretraining.

### v2

@@ 4.4 Downstream Task: ECG-based User Identification
ECG generally shows unique patterns that can distinguish individuals, which makes them suitable for subject identification tasks and offers potential advantages over other biometric traits (Melzi et al., 2023). For example, compared to identifying persons with facial images, using ECG can further protect users’ privacy during usage. In this study, we leverage the PTB-XL (Wagner et al., 2020) and ICBEB (Liu et al., 2018) datasets, also used in the arrhythmia diagnosis task, to design a one-shot learning benchmark for ECG identification. Similar to the arrhythmia diagnosis task in Section 4.3, we focus solely on the test splits of the PTB-XL and ICBEB datasets. For PTB-XL, we select 5-second signal sequences from each of the 1907 subjects for both training and testing, which results in a 1907-class classification task. Similarly, we select 4-second samples and classes from the ICBEB dataset, which forms a 689-class classification task. In this task, we compare the performance of the proposed ESI method with various approaches under linear probing and fine-tuning settings. For the specialized supervised methods, we compared our method with established supervised learners, including LSTM, XResNet101, ResNet50, and ensemble methods (Strodthoff et al., 2020). Besides the supervised learning methods, we also evaluated the ESI method against signal-focused SSL approaches including SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), CLOCS (Kiyasseh et al., 2021), and LEAVES (Yu et al., 2022a) under the linear probing and fine-tuning settings. Due to the patient de-identification during pretraining, we cannot perform zero-shot learning for this identification task, as the model has not been exposed to identifiable patient information.
@@
Table 4: Evaluation results of 1-shot ECG-based user identification task under settings including supervised learning, linear probing, and fine-tuning. The metric used is the area under the ROC curve (AUC) and accuracy score (ACC). The best results are highlighted in bold.
@@
LSTM (Strodthoff et al., 2020) 0.907±0.014 0.444±0.012 0.918±0.011 0.610±0.011
XResNet101 (Strodthoff et al., 2020) 0.915±0.011 0.473±0.015 0.921±0.017 0.623±0.020
ResNet50 (Strodthoff et al., 2020) 0.926±0.016 0.497±0.021 0.933±0.013 0.641±0.015
Ensemble (Strodthoff et al., 2020) 0.930±0.011 0.500±0.015 0.937±0.013 0.653±0.016
ConvNeXt-Tiny (Woo et al., 2023) 0.918±0.014 0.480±0.018 0.936±0.014 0.647±0.016
ConvNeXt-Base (Woo et al., 2023) 0.922±0.013 0.494±0.021 0.932±0.013 0.649±0.015
@@
SimCLR (Chen et al., 2020) 0.806±0.022 0.185±0.033 0.838±0.019 0.252±0.021
BYOL (Grill et al., 2020) 0.837±0.019 0.240±0.027 0.855±0.019 0.283±0.035
CLOCS (Kiyasseh et al., 2021) 0.822±0.022 0.207±0.036 0.841±0.020 0.235±0.037
LEAVES (Yu et al., 2022a) 0.840±0.017 0.248±0.029 0.853±0.022 0.275±0.029
(Ours) ESI-tiny 0.923±0.014 0.510±0.017 0.937±0.012 0.654±0.016
(Ours) ESI 0.927±0.011 0.517±0.014 0.944±0.010 0.665±0.012
@@
SimCLR (Chen et al., 2020) 0.936±0.011 0.522±0.014 0.945±0.012 0.673±0.015
BYOL (Grill et al., 2020) 0.942±0.011 0.547±0.014 0.951±0.013 0.686±0.012
CLOCS (Kiyasseh et al., 2021) 0.929±0.009 0.516±0.010 0.940±0.011 0.666±0.010
LEAVES (Yu et al., 2022a) 0.944±0.011 0.550±0.013 0.953±0.009 0.688±0.015
(Ours) ESI-tiny 0.966±0.010 0.591±0.014 0.980±0.007 0.747±0.012
(Ours) ESI 0.970±0.009 0.608±0.013 0.985±0.008 0.762±0.010
@@ 5 Discussion
Table 6: Ablation experiment on pretraining ESI-Tiny on the MIMIC-IV-ECG data only. The table compares the performance of the model pre-trained exclusively on MIMIC-IV-ECG with the model pre-trained on the full dataset across different tasks and datasets.
Pretraining Set Arrhythmia (Linear Probing) Arrhythmia (Zero-shot) Identification (Linear Probing)
PTB-XL ICBEB PTB-XL ICBEB PTB-XL ICBEB
MIMIC 0.912 0.967 0.792 0.832 0.917 0.934
All 0.928 0.976 0.807 0.842 0.923 0.937
diff $\downarrow$ 0.016 $\downarrow$0.009 $\downarrow$0.015 $\downarrow$ 0.010 $\downarrow$ 0.006 $\downarrow$ 0.003
@@ 5.1 Ablation Study: Component Analysis and Impact
5.2 Ablation Study: Impact of Pre-Training with MIMIC-IV-ECG Only
As the MIMIC-IV-ECG data is the single dataset that contains the most ECG samples used in pretraining our ESI models, to investigate the impact of using only the MIMIC-IV-ECG dataset for pre-training, we conducted an experiment where the model was pre-trained exclusively on MIMIC-IV-ECG. We then evaluated the model’s performance on the PTB-XL and ICBEB datasets, focusing on three tasks: arrhythmia detection (both linear probing and zero-shot) and subject identification (linear probing).
Table 6 presents the performance results across these tasks. The results indicate that pre-training on the full dataset yields better performance across all tasks and datasets compared to pre-training exclusively on MIMIC-IV-ECG. Specifically, there is a noticeable improvement in both arrhythmia detection and subject identification tasks when additional datasets are included in the pre-training process. For the arrhythmia detection task, the difference in performance is more substantial in the zero-shot setting, where the model pre-trained on all datasets achieves an improvement of 0.015 in PTB-XL and 0.014 in ICBEB. In the identification task, the model pre-trained on all datasets only shows a slight improvement.
These results suggest that while MIMIC-IV-ECG provides a strong foundation for pre-training, and the inclusion of additional datasets can further enhance the model’s ability to generalize across different tasks and datasets. Also, for the zero-shot learning tasks, learning from the samples and annotating information from the same dataset during pretraining, e.g., PTB-XL, could help the downstream tasks accordingly.
@@ 5.3 Ablation Study: Selection of Backbone for Signal Encoder
Table 5b(b) summarizes the results. ConvNeXt-based models (ConvNeXt-tiny and ConvNeXt-base) achieved the highest AUC scores on both arrhythmia diagnosis and ECG identification tasks. Notably, ConvNeXt-base, the largest model with the most parameters (85.56M), provides the best overall performance (AUC of 0.932 for arrhythmia diagnosis and 0.926 for ECG identification). ConvNeXt-tiny, a more parameter-efficient option (26.81M), shows a slightly lower performance compared to the base version, but also achieves competitive AUC scores with a significantly lower parameter footprint. With the same-level size as ConvNeXt models in both tasks, the ViT model shows substantially lower performances, which might indicate the convolutional kernel could be more suitable for processing ECG signals. Moreover, the XResNet1D101 backbone displays lower AUC scores compared to ConvNeXt architectures. This can be due to the significantly smaller parameter size of the XResNet model.
@@ 5.4 Ablation Study: Impact of Pretraining Data Sizes and Distribution Shifts
In this ablation study, we also investigate the impact of pretraining dataset size on the performance of our ESI-tiny model in two arrhythmia classification datasets, including PTB-XL and ICBEB. We artificially change the size of the unlabeled pretraining dataset from 0 to all available samples by randomly sampling and selecting from all the pretraining samples. After ESI-tiny encoders are trained, the models are then evaluated on both datasets using a linear probing approach with frozen encoder weights, and the AUC score is employed as the performance metric.
@@ 6 Conclusion
This study introduces a novel multimodal contrastive pretraining framework to enhance the quality and robustness of representations learned from ECG signals. To address the lack of descriptive text associated with ECGs, we propose a retrieval-augmented generation (RAG) pipeline called the Cardio Query Assistant (CQA). This pipeline generates detailed textual descriptions for ECG data with demographic information, potential conditions, and waveform patterns. Inspired by the success of multimodal pretraining strategies in vision-language tasks, we develop the ECG Semantic Integrator (ESI). This framework integrates both contrastive and captioning capabilities to foster a deeper semantic understanding of ECG signals. Our evaluation validates the effectiveness of the proposed approach across several downstream tasks. The ESI method demonstrates improvement in arrhythmia diagnosis and ECG-based user identification tasks by outperforming strong baselines that cover supervised learning and SSL approaches. These results with the ablation studies highlight the benefits of multimodal learning for ECG analysis and the value of integrating captioning loss with contrastive pretraining. Beyond ECG, we believe the proposed CQA and ESI frameworks hold the potential for applications to other types of biomedical time series data, where contextual information can be leveraged to enhance the representation learning and downstream analysis.
@@
Acknowledgments
This work was supported by National Science Foundation (# 2047296) and National Institute of Health (# R01DA059925)
Appendix A Ablation Study on CQA
The introduction of Retrieval-Augmented Generation (RAG) in CQA in this study plays a crucial role in enriching the pretraining dataset by generating detailed ECG waveform descriptions. In this ablation study, we investigate the impact of different RAG settings on performance.
To evaluate the performance of our RAG-generated descriptions, we utilize the human-annotated ECG waveform descriptions from (Wen & Kang, 2022) as references. This dataset includes annotations for 28 types of arrhythmia. We generate corresponding descriptions using our RAG method and compared them against the reference descriptions using BERTScore (Zhang et al., 2020) with the RoBERTa-large model (Liu et al., 2019). BERTScore is a metric designed to evaluate the similarity between generated text and reference text by leveraging pre-trained contextual embeddings from transformer models. BERTScore computes the similarity of embeddings to provide a more nuanced assessment of semantic alignment, and there are three main metrics served in BERTScore:
  • Precision measures how well the tokens in the generated text match with the tokens in the reference text, which indicates how accurately the generated content reflects the reference.
  • Recall assesses how well the tokens in the reference text are captured by the generated text, which reflects the completeness of the generated content.
  • F1 Score is the harmonic mean of Precision and Recall, which provides a balanced measure of both accuracy and completeness in the semantic alignment between the generated and reference texts.
A.1 Impact of Top-k in Vector Search
In this section, we examine the effect of varying the top-k parameter in the vector search interface on the model’s performance. The top-k parameter controls how many of the top-ranked candidates are considered in the generation process. We evaluated the model’s performance by measuring Precision, Recall, and F1 score generated by BERTScore for different values of top-k. The mean and standard deviation of these metrics across multiple runs are summarized in Table 7.
Table 7: BERTScores across different Top-k values
Top K $\textbf{B}_{Precision}$uparrow $\textbf{B}_{Recall}$uparrow $\textbf{B}_{F1Score}$uparrow
1 0.8190 (0.0128) 0.8326 (0.0237) 0.8256 (0.0147)
2 0.8197 (0.0159) 0.8402 (0.0156) 0.8296 (0.0090)
3 0.8169 (0.0160) 0.8433 (0.0155) 0.8297 (0.0106)
4 0.8159 (0.0157) 0.8447 (0.0160) 0.8299 (0.0111)
5 0.8153 (0.0172) 0.8447 (0.0167) 0.8295 (0.0110)
The results indicate that as the top-k value increases, the Recall generally improves slightly, which suggests that considering more candidates helps in retrieving more relevant information. However, the mean Precision tends to decrease marginally, which reflects a trade-off where including more candidates might introduce less relevant information. The F1 score, which balances both Precision and Recall, remains relatively stable across different top-k values. This indicates that the model maintains a consistent performance overall.
Our findings suggest that varying the top-k parameter has a limited but consistent impact on RAG performance. The choice of top-k can be adjusted based on the specific needs of the task, with higher values favoring recall at a slight cost to precision, while lower values provide a more precise but slightly less comprehensive output. In this study, we set the top-k value as 2.
A.2 Impact of the Number of Textbooks in RAG
We investigate the impact of the number of textbooks used as the knowledge base in the CQA pipeline. The purpose of this ablation study is to assess whether the inclusion of more or fewer textbooks affects the quality of the generated ECG waveform descriptions. The textbooks used in this study are "ECG Workout: Exercises In Arrhythmia Interpretation" by Huff (2006) and "12-Lead ECG: The Art of Interpretation" by Garcia (2015). We conducted experiments using three different configurations: using no textbooks, using each textbook individually, and using both textbooks combined. The results are summarized in Table 8.
Table 8: BERTScores across different numbers of textbooks used in RAG
Textbooks $\textbf{B}_{Precision}$uparrow $\textbf{B}_{Recall}$uparrow $\textbf{B}_{F1Score}$uparrow
No Textbook 0.7617 (0.0187) 0.8154 (0.0142) 0.7928 (0.0126)
Garcia (2015) 0.8169 (0.0142) 0.8290 (0.0175) 0.8217 (0.0111)
Huff (2006) 0.8182 (0.0145) 0.8395 (0.0170) 0.8285 (0.0100)
Both 0.8197 (0.0159) 0.8402 (0.0156) 0.8296 (0.0090)
The results show a clear trend that incorporating textbooks into the RAG pipeline improves the performance across all BERTScore metrics (Precision, Recall, and F1 Score). Specifically, using both textbooks in combination yields the highest scores, with a Precision of 0.8197, a Recall of 0.8402, and an F1 score of 0.8296. This suggests that the inclusion of multiple sources of knowledge enriches the generated descriptions. When using only one textbook, Huff (2006) slightly outperforms Garcia (2015) in terms of all three metrics, although the difference is incremental. This indicates that while each textbook independently contributes to the generation quality, combining them provides a more robust knowledge base for the RAG process. In contrast, the absence of textbooks results in a significant drop in performance, with the lowest scores across all metrics. This highlights the importance of incorporating domain-specific knowledge to improve the relevance and accuracy of LLMs in generating content. Thus, in this study, using multiple textbooks in the RAG pipeline not only increases the quality of the generated ECG descriptions but also ensures consistency in the output.
A.3 Impact of Different LLM Prompting Strategies
In this subsection, we investigate the effect of varying the prompt structure used in the LLM within the RAG pipeline on the quality of the generated ECG waveform descriptions. To assess consistency of the generation across different prompts, we conducted an ablation study by rephrasing our standard prompt, "How is {ECG condition} reflected in 12-lead ECG?", into five different variations such as "Describe how {ECG condition} appears in a 12-lead ECG." and "What are the ECG findings associated with {ECG condition} in a 12-lead setup?" The alternative prompts used are designed to explore whether slight changes in wording would affect the performance of the RAG pipeline. Despite these variations, the underlying query remains focused on generating descriptions of how specific ECG conditions are in a 12-lead ECG.
The results demonstrate a remarkable consistency across all rephrased prompting strategies, with an average mean F1 score of 0.8295 and a very low standard deviation of 0.0003 across the variations of prompts. This indicates that the CQA’s ability to generate relevant descriptions is robust to variations in the phrasing of the prompt. This robustness allows for some flexibility in how prompts are formulated without hurting the quality of the generated content.
Appendix B Datasheet for ECG Datasets Used for Training ECG Model
In this section, we compile a dataset for the used training data following Gebru et al. (2021).
B.1 Motivation
Purpose: The datasets were collected and processed to facilitate the development and evaluation of automated ECG processing algorithms, particularly for training a foundation model that extracts generalized embedding from ECG signals. These datasets provide a large number of ECG recordings paired with human or machine annotations and metadata, which enables robust training of machine learning models in the medical domain.
Creators: The datasets were created by institutions including PhysioNet, Chapman University, Shaoxing People’s Hospital, and Physikalisch-Technische Bundesanstalt (PTB).
Funding: Development and release of these datasets were supported by various grants, including the NIH, the Kay Family Foundation, and the Berlin Big Data Center.
B.2 Composition
Instances:
  • MIMIC-IV-ECG: Contains diagnostic 12-lead ECG recordings, with each instance representing an ECG signal paired with machine-generated annotations and metadata.
  • PTB-XL: Comprises 17,415 clinical 12-lead ECG records, annotated by cardiologists, to cover a wide range of diagnostic categories.
  • Chapman-Shaoxing: Includes 45,152 12-lead ECGs from patients, annotated by professional experts, with conditions such as atrial fibrillation, myocardial infarction, and other arrhythmias.
Volume:
  • MIMIC-IV-ECG: The dataset includes thousands of ECG records.
  • PTB-XL: Contains 17,415 ECG records.
  • Chapman-Shaoxing: Contains 45,152 ECG records.
Data Features: Each dataset includes ECG waveform data, demographic information (age, sex, etc.), and clinical annotations related to cardiac conditions. PTB-XL and Chapman-Shaoxing datasets also include metadata on the diagnostic process, such as SCP-ECG statements.
B.3 Collection Process
Data Acquisition:
  • MIMIC-IV-ECG: Data was collected as part of routine clinical care and stored in the MIMIC-IV database.
  • PTB-XL: ECG signals were collected using devices from Schiller AG over a period of seven years, with annotations provided by cardiologists.
  • Chapman-Shaoxing: Data was collected from multiple hospitals, with ECG signals recorded and labeled by licensed physicians.
Ethical Considerations: All datasets are publicly available and de-identified to protect patient privacy. The collection of the original data was approved by relevant institutional ethics committees. No new data was collected in this study.
B.4 Preprocessing/Cleaning/Labeling
MIMIC-IV-ECG: Machine-generated annotations were used for labeling. The data was cleaned to remove noise and artifacts.
PTB-XL: ECG data was preprocessed to ensure consistency in the waveform format, and SCP-ECG codes were applied for standardized labeling.
Chapman-Shaoxing: ECG data was denoised, and diagnostic information was verified by multiple physicians before being stored in a structured format.
B.5 Uses
Current Uses: These datasets have been used for developing and evaluating machine learning models for ECG analysis, particularly in pretraining ECG models and detecting arrhythmias.
Potential Uses: The datasets can be used for broader research in cardiac health, including developing diagnostic tools and studying the epidemiology of heart conditions.
Restrictions: While the datasets are de-identified, users should still adhere to ethical guidelines when using the data for research.
B.6 Distribution
Access: The datasets are publicly available through the PhysioNet platform under the Creative Commons Attribution 4.0 International License.
Regulatory Issues: There are no specific regulatory restrictions, but users must follow the licensing terms.
B.7 Maintenance
Support: Users can contact the dataset maintainers through the PhysioNet platform for support or questions.
Updates: Any updates or new versions of the datasets will be announced on the PhysioNet platform.
Community Contributions: Researchers are encouraged to contribute their findings and improvements back to the community through publications or direct contributions to the dataset repository.